### v2

# Score-based Conditional Out-of-Distribution Augmentation for Graph Covariate Shift

Abstract.
Distribution shifts between training and testing datasets significantly impair the model performance on graph learning. A commonly-taken causal view in graph invariant learning suggests that stable predictive features of graphs are causally associated with labels, whereas varying environmental features lead to distribution shifts. In particular, covariate shifts caused by unseen environments in test graphs underscore the critical need for out-of-distribution (OOD) generalization. Existing graph augmentation methods designed to address the covariate shift often disentangle the stable and environmental features in the input space, and selectively perturb or mixup the environmental features. However, such perturbation-based methods heavily rely on an accurate separation of stable and environmental features, and their exploration ability is confined to existing environmental features in the training distribution. To overcome these limitations, we introduce a novel distributional augmentation approach enabled by a tailored score-based conditional graph generation strategies to explore and synthesize unseen environments while preserving the validity and stable features of overall graph patterns. Our comprehensive empirical evaluations demonstrate the enhanced effectiveness of our method in improving graph OOD generalization.
$c$s: Computing methodologies Neural networksccs: Mathematics of computing Graph algorithms
1. Introduction
Deep learning has become the dominant paradigm for analyzing graph-structured data across diverse domains. However, most graph learning methods assume that training and test graphs are independently and identically distributed (i.i.d.)., an assumption that frequently fail in real-world setting (gui2022good; huang2024enhancing). On the Web, spanning hyperlink graphs, social networks, recommendation and e-commerce interaction graphs, and knowledge graphs, distribution shift frequently arise due to product launches, policy or UI changes, seasonality, bot activity, and rapid community growth. Prior studies have shown that such shifts can lead to substantial performance degradation (gui2022good; sui2024unleashing), motivating growing interest in graph out-of-distribution (OOD) generalization, including invariant graph learning (chen2022learning; wu2022discovering; huang2024enhancing) and graph data augmentation (rong2019dropedge; wang2021mixup; han2022g; yao2022improving; sui2024unleashing; ligraph).
Following (gui2022good; sui2024unleashing), we distinguish correlation shift, where environment–label correlations differ but test environments are represented during training, from covariate shift, where the test environments are not present in the training dataset. A causal view in invariant graph learning (chen2022learning; wu2022discovering) posits that label-determining stable patterns are invariant across environments, while environmental patterns vary. Prior work largely targets correlation shift by isolating stable patterns (miao2022interpretable; chen2022learning; wu2022discovering). We focus instead on the underexplored covariate shift in graph learning, which is pervasive in Web settings (e.g., emerging communities or new interaction regimes unseen at train time).
Existing graph augmentation techniques typically perturb graph structures or features by mixing or dropping edges (rong2019dropedge; han2022g; sui2024unleashing). While effective with regularized contexts, such operations offer limited exploration ability to unseen environments and risk generating semantically inconsistent or unreliable samples (e.g., invalid molecules in chemistry), as indiscriminate edits may distort label-relevant stable patterns, yielding uncontrolled augmented distributions. Controlled methods (sui2024unleashing) help but often require an explicit, accurate split of stable vs. environmental subgraphs—nontrivial and sometimes infeasible on Web graphs. This raises a key question:
Can we explore beyond the training environments in a controllable way while preserving label-critical stable patterns?
We provide an affirmative answer via a score-based conditional graph generation strategy for covariate-shift mitigation. We formulate OOD augmentation as generating graphs conditioned on a target label and an exploration parameter that controls the degree of distributional deviation, building on the widely used graph-generation hypothesis (wu2022discovering; yang2022learning; gui2022good; chen2022learning). Inspired by diffusion models (song2020score), we define a forward corruption process and a novel reverse denoising process that jointly (i) retain predictive stable patterns and (ii) inject environment exploration. Our conditional score function incorporates both label information and the degree of OOD-ness, enabling Conditional Out-of-Distribution Diffusion Augmentation (CODA) to generate graphs likely to preserve stable patterns while probing underrepresented (low-density) regions [^1] Low-density regions are parts of the class-conditional data manifold with few samples; see Section 4.1. of the class-conditioned manifold beyond the training graph space.
CODA inherits diffusion-model robustness, helping ensure validity (preventing, e.g., non-viable molecules) without explicitly separating stable from environmental subgraphs. We evaluate on synthetic and real-world graph classification tasks under diverse covariate shifts, which are commonly observed in Web contexts. For instance, motif-frequency variations emulate evolving community or interaction structures; changes in graph size correspond to thread-depth or document-length drift; and feature/structural shifts reflect the dynamics of user/item attributes and connectivity patterns in online platforms. Across these settings, CODA achieves consistent gains over state-of-the-art invariant-learning and augmentation methods, enabling controlled exploration of OOD environments while preserving stable patterns.
Our main contributions are summarized as follows:
  • We propose a novel conditional graph generation-based augmentation framework that addresses covariate distribution shifts. Our method enables controlled exploration of environmental patterns while preserving stable patterns, without requiring explicit subgraph decomposition.
  • We develop a unified generator for OOD structures, node features, and edge features, handling feature, structural, and joint shifts which are pervasive in Web contexts.
  • Extensive experiments across synthetic, semi-artificial, molecular, and Web-analogous text/review graph datasets demonstrate consistent gains over strong OOD generalization baselines.
2. Related Work
Graph-structured data, such as hyperlink graphs, social networks and e-commerce interaction graphs in Web, are inherently complex, characterized by the intricate challenges of irregularity and nuanced structural information. This complexity gives rise to graph OOD problems that not only necessitate addressing shifts in feature distributions but also demand attention to variations of structural distributions. In this context, we summarize two principal categories of algorithms for graph OOD robustness: (i) invariant graph learning strategies, which aim to ensure model stability across varying distributions; and (ii) graph data augmentation techniques, designed to enhance model generalizability by simulating diverse distribution scenarios.
Invariant Graph Learning. The concept of invariant graph learning, inspired from seminal works (arjovsky2019invariant; rosenfeld2020risks; ahuja2021invariance), targets on identifying stable graph structures (e.g., subgraphs) or representations (predictors) that remain consistent across different environments, thereby enhancing OOD generalization. This is achieved by capturing salient graph features and minimizing empirical risks across varying conditions. In scenarios where establishing causality is complex or where strong assumptions may not hold, the task can be approximated by identifying features that demonstrate invariance under distributional shifts, thereby facilitating OOD generalization (li2022out). Effective OOD generalization is achieved by basing predictions solely on invariant information (li2022out). For example, DIR (wu2022discovering) distinguishes between invariant and environment-specific subgraphs by creating varied interventional distributions on the training distribution. CIGA (chen2022learning) further explores this domain by employing synthetic environments and the graph generation process to identify stable features under various distribution shifts. However, this line of research assumes access to test environments during training, which is an unrealistic assumption given the impracticality of covering all possible test scenarios. Training in limited environments reduces spurious correlations but fails to generalize to new, unseen environments. DISGEN (huang2024enhancing) gains promising results in disentangling the size factors from graph representations by minimizing the shared information between size- and task-related information, however, the technique is constrained to handle size generalization. Recently, CANET (wu2024graph) introduces an environment estimator that infers a pseudo-environment, enabling the model to capture stable features under node-level distribution shifts without requiring prior knowledge of environment labels. Nevertheless, CANET does not address OOD issues in other applications, such as molecular graphs. In this work, we propose a framework capable of generalizing to unseen environments characterized by differences not only in graph size but also in graph structure, node features, and edge features.
Graph Data Augmentation. Beyond invariant graph learning, graph data augmentation aims to diversify the training distribution, thereby enhancing the OOD generalization of models. DropEdge (rong2019dropedge) introduces randomness by selectively removing edges, thus varying the training data’s structure. M-Mixup (wang2021mixup) enriches the dataset by interpolating diverse and irregular graphs within semantic space. $\mathcal{G}$-Mixup (han2022g) extends mixup to graph classification by interpolating across graph generators (graphons). Adversarial augmentation techniques, such as FLAG (kong2022robust), perturb node features via gradients, while AIA (sui2024unleashing) learns adversarial masks to probe environmental discrepancies. Despite these advances, methods that edit only the observed training graphs often have limited ability to explore truly novel environments and can distort label-determining patterns, yielding uncontrolled augmentations. Environment-aware frameworks (ligraph) partially address this by using environment metadata to linearly explore structures and features; however, they rely on high-quality, diverse environment labels that are costly and often infeasible to obtain. GRATIN (abbahaddou2025graph) takes a different route: it fits a Gaussian mixture on learned graph embeddings and samples synthetic representations—supported by a Rademacher-complexity analysis—to achieve efficient gains in generalization without editing raw graphs. In contrast, our work introduces a generation-based augmentation method that dispenses with environment labels and generated graphs end-to-end (structures, node features, and edge features), enabling controlled exploration beyond the training graph space while preserving stable, label-determining patterns.
3. Problem Formulation
Notations. We represent a graph with $n$ nodes as $G=(\boldsymbol{A},\boldsymbol{X},\boldsymbol{E})$, where $\boldsymbol{A}\in\mathbb{R}^{n\times n}$ is the adjacency matrix, $\boldsymbol{X}\in\mathbb{R}^{n\times a}$ denotes $a$-dimensional node features and $\boldsymbol{E}\in\mathbb{R}^{n\times n\times b}$ encodes $b$-dimensional edge features. Without the loss of generality, we focus on the graph classification task where each graph $G$ is associated with a label $Y\in\mathcal{Y}$, determined by a predefined labelling rule $\mathcal{G}\rightarrow\mathcal{Y}$. Following invariant learning (ahuja2021invariance; chen2022learning), we denote the graph dataset as $\mathcal{D}=\{(G_{i}^{e},Y_{i}^{e})\}_{e\in\mathcal{E}_{\mathrm{all}}}$, where $(G_{i}^{e},Y_{i}^{e})\sim P_{e}(G,Y)$ is an i.i.d. draw in the environment $e$ sampled from all possible environments $\mathcal{E}_{\mathrm{all}}$. The complete dataset can be partitioned into a training set $\mathcal{D}_{\mathrm{tr}}=\{(G_{i}^{e},Y_{i}^{e})\}_{e\in\mathcal{E}_{\mathrm{tr}}}$ and a test set $\mathcal{D}_{\mathrm{te}}=\{(G_{i}^{e},Y_{i}^{e})\}_{e\in\mathcal{E}_{\mathrm{te}}}$, where $\mathcal{E}_{\mathrm{tr}}$ and $\mathcal{E}_{\mathrm{te}}$ index the training and testing environments, respectively. In practice, environment information may not be explicitly given, and we further denote the training distribution as $P_{\mathrm{tr}}(G,Y)$ and the testing distribution as $P_{\mathrm{te}}(G,Y)$.
3.1. Graph Covariate Shift
We study OOD graph classification under covariate shift: the label mechanism is stable while the input distribution shifts, i.e.,
$P_{\mathrm{tr}}(Y\mid G)=P_{\mathrm{te}}(Y\mid G)\quad\text{but}\quad P_{\mathrm{tr}}(G)\neq P_{\mathrm{te}}(G).$
This mismatch arises from limited training coverage and changing test conditions. Prior work (ligraph) distinguishes: (i) feature shift $P_{\mathrm{tr}}(\mathbf{X})\neq P_{\mathrm{te}}(\mathbf{X})$ with $P_{\mathrm{tr}}(\mathbf{A},\mathbf{E})=P_{\mathrm{te}}(\mathbf{A},\mathbf{E})$ (e.g., GOOD-CMNIST (gui2022good), where color defines environments); (ii) structural shift $P_{\mathrm{tr}}(\mathbf{A},\mathbf{E})\neq P_{\mathrm{te}}(\mathbf{A},\mathbf{E})$; and (iii) joint shift $P_{\mathrm{tr}}(\mathbf{A},\mathbf{X},\mathbf{E})\neq P_{\mathrm{te}}(\mathbf{A},\mathbf{X},\mathbf{E})$ (e.g., GOOD-Motif size/basis changes (gui2022good), or scaffold splits in GOOD-HIV (gui2022good)). All three forms challenge OOD generalization; our framework is designed to address them.
3.2. Graph Classification under Covariate Shift
With only observing the training set $\mathcal{D}_{\mathrm{tr}}$ sampled from the training distribution $P_{\mathrm{tr}}$ in training environments $\mathcal{E}_{\mathrm{tr}}$, our generalization objective under graph covariate shift is to train an optimal graph classifier $f:\mathcal{G}\rightarrow\mathcal{Y}$ that performs well across any possible environments $\mathcal{E}_{\mathrm{all}}\supseteq\mathcal{E}_{\mathrm{tr}}$. We formulate this goal as the following minimization problem:
(1) $\underset{f}{\min}\ \mathbb{E}_{e\in\mathcal{E}_{\mathrm{all}}}\mathbb{E}_{(G^{e},Y^{e})\sim P_{e}(G,Y)}[\ell(f(G^{e}),Y^{e})],$
where $\ell(\cdot,\cdot)$ denotes the loss function for graph classification and the expectation is with respect to graphs under all possible environments. However in practice, the training environments $\mathcal{E}_{\mathrm{tr}}$ may not cover all environments, causing degraded classification performance when applying the learned classifier in unseen test environments. This covariate shift calls for an effective manner to sufficiently explore unseen data distribution or environments during model training. We summarize and discuss the various types of graph covariate shifts in detail in Appendix.
3.3. Issues in Environmental Augmentation
To augment the training distribution for mitigating graph covariate shift, existing solutions often approach Eq. (1) by separating and augmenting the environments:
(2) ${\min_{f}}\ \mathbb{E}_{e\in\{\mathcal{E}_{\mathrm{tr}}\cup\mathcal{E}_{\mathrm{aug}}\}}\mathbb{E}_{(G^{e},Y^{e})\sim P_{e}(G,Y)}[\ell(f(G^{e}),Y^{e})],$
where the augmented environments $\mathcal{E}_{\mathrm{aug}}$ are obtained based on either interpolating explicitly given environmental labels (ligraph) or perturbing implicitly separated environmental components (sui2024unleashing; wu2022discovering; miao2022interpretable; chen2022learning). Obtaining accurate environmental labels and components itself could be high-cost and nontrivial tasks. Furthermore, without explicit environmental information and additional assumptions, separating environmental components could be inherently unfeasible (chen2024does), which limit the practicability of such strategy. Additionally, the subgraph perturbations based augmentation is mainly operated by edge dropping (sui2024unleashing; rong2019dropedge) and mixup (han2022g), which is confined to existing subgraphs in training data and could cause invalid samples (e.g., generating molecules that are chemically invalid).
4. Score-based Conditional Out-of-Distribution Diffusion Augmentation (CODA)
In this section we present CODA, a score-based augmentation framework that generates graph samples exploring new environments while preserving label-determining structure. We first introduce a distribution-level view of augmentation, then describe the forward and reverse diffusion processes, and finally formalize the working principles of CODA.
Our New Perspective: Distributional Augmentation with OOD Control. To overcome the limitations of existing strategies, this work formulates the augmentation problem as a generation-based graph OOD augmentation strategy, which directly models and augments the training distribution, without explicitly requiring the knowledge or separation of environmental information. Specifically, we target on synthesizing an augmented training distribution $\tilde{P}_{\mathrm{tr}}(G,Y)$, which is combined with the original training distribution to obtain the classifier, stated as:
(3) $\underset{f}{\min}\ \mathbb{E}_{(G,Y)\sim\{P_{\mathrm{tr}}(G,Y)\cup\tilde{P}_{\mathrm{tr}}(G,Y)\}}[\ell(f(G),Y)].$
The augmented distribution $\tilde{P}_{\mathrm{tr}}(G,Y)$ can be implemented in multiple ways, but it needs to satisfy two principles: (i) $\tilde{P}_{\mathrm{tr}}(G,Y)$ should deviate from $P_{\mathrm{tr}}(G,Y)$ in a controlled manner for exploration, and (ii) the explored graphs in $\tilde{P}_{\mathrm{tr}}(G,Y)$ should preserve the stable patterns of graphs in $P_{\mathrm{tr}}(G,Y)$. However, current graph generation models (jo2022score; martinkus2022spectre; vignac2022digress) cannot directly generate graphs that meet these two criteria. To address this, we propose a novel score-based conditional generative model that captures the augmented distribution $\tilde{P}_{\mathrm{tr}}(G,Y)$ while adhering to both principles.
[…]
(4) $P_{\mathrm{tr}}\left(G,Y\mid\mathcal{E}_{\text{ood }}=\lambda\right)\propto p\left(G\right)p\left(Y\mid G\right)p\left(\mathcal{E}_{\text{ood }}=\lambda\mid G,Y\right).$
[…]
4.1. Preliminaries
Causal View. Following (wu2022discovering), we assume that each input graph $G$ can be decomposed into a causally relevant component $C$ and a spurious (environment-specific) component $S$, e.g., the House motif vs. the Tree base in GOOD-Motif (gui2022good). The label $Y$ depends only on $C$, while $S$ captures variations across environments. These qualitative causal assumptions (pearl2016causal; neuberg2003causality) can be summarized as:
  • $C\to G\leftarrow S$  ($G$ consists of $C$ and $S$).
  • $C\to Y$  ($C$ is sufficient to determine $Y$).
  • $C\not\!\perp\!\!\!\perp S$  ($C$ and $S$ may be statistically dependent, e.g., via latent confounding).
  • $S\perp\!\!\!\perp Y\mid C$  (given $C$, $S$ carries no more information about $Y$).
Low-Density Regions. Class-conditional densities are typically long-tailed (sehwag2022generating), with low-density neighborhoods on the data manifold containing few or no training samples. OOD graphs $G_{\text{ood}}$ tend to reside in these regions (lee2023exploring), and diffusion models are observed to interpolate rather than memorize in such regions (sehwag2022generating). In CODA, we introduce an exploration parameter $\lambda$ to control the extent to which generation is encouraged toward these sparse regions.
Let $\tilde{P}_{e}$ denote the generator’s joint distribution at exploration level $e\equiv\mathcal{E}_{\text{ood}}(\lambda)$. When exploration is disabled ($e=0$), the generator should reproduce the in-distribution data:
$\tilde{P}_{0}(G,Y)\approx P_{\text{in}}(G,Y).$
As the exploration level increases, the generator should allocate more probability mass to OOD neighborhoods of the data manifold. Writing $\mathcal{G}_{\text{ood}}$ for a measurable subset of low-density OOD regions and taking $e_{2}>e_{1}\geq 0$, we expect
$\tilde{P}_{e_{2}}\!\big(G\in\mathcal{G}_{\text{ood}},Y\big)\;>\;\tilde{P}_{e_{1}}\!\big(G\in\mathcal{G}_{\text{ood}},Y\big),$
that is, in shorthand,
$\tilde{P}(G_{\text{ood}},Y\mid\mathcal{E}_{\text{ood}}>0)\;>\;\tilde{P}(G_{\text{ood}},Y\mid\mathcal{E}_{\text{ood}}=0).$
4.2. Forward Diffusion Process
In this work, we focus on diffusion over continuous features to address distribution shifts that occur not only in discrete graph structures but also in continuous node and edge features. The foundational work (song2020score) introduced a method for modeling the diffusion process of data into noise and vice versa using stochastic differential equations (SDEs). For graph generation, this diffusion process gradually corrupts graphs into a prior distribution like the normal distribution. As illustrated in Figure 1, given an unlabeled graph $G$, we use continuous time $t\in[0,T]$ to index the diffusion steps $\left\{G_{t}\right\}_{t=1}^{T}$ of the graph, where $G_{0}$ represents the original distribution and $G_{T}$ follows a prior distribution. The forward diffusion process from the graph to the prior distribution is defined via an Itô SDE:
(5) $\mathrm{d}G_{t}=\mathbf{f}_{t}\left(G_{t}\right)\mathrm{d}t+g_{t}\mathrm{~d}\mathbf{w},$
which incorporates linear drift coefficient $\mathbf{f}_{t}(\cdot):\mathcal{G}\rightarrow\mathcal{G}$ and scalar diffusion coefficient $g_{t}:\mathcal{G}\rightarrow\mathbb{R}$ related to the amount of noise corrupting the unlabelled graph at each infinitesimal step $t$, along with a standard Wiener process $\mathbf{w}$.
Figure 1. The diffusion and reverse processes of OODA. The diffusion process iteratively transforms an unlabeled original graph $G_{0}$ into noise $G_{T}$. During the denoising process, $\tilde{G}_{t-1}$ is computed using the conditional score $\nabla_{G_{t}}\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)$, constrained by the target class $\mathbf{y}_{G}$ and the exploration parameter $\lambda$. Ultimately, the clean OOD graph $\tilde{G}_{0}$ is generated.
4.3. Score-based Conditional Graph Generation with OOD Control
To steer generation toward controlled OOD characteristics, we condition the reverse-time SDE on the target class $\mathbf{y}_{G}$ and an exploration knob $\lambda\in[0,1]$ that modulates OOD-ness:
(6) $\mathrm{d}G_{t}=\Big[\mathbf{f}_{t}(G_{t})-g_{t}^{2}\,\nabla_{G_{t}}\log p_{t}\!\big(G_{t},\mathbf{y}_{G}\mid\mathcal{E}_{\mathrm{ood}}=\lambda\big)\Big]\mathrm{d}\bar{t}\;+\;g_{t}\,\mathrm{d}\overline{\mathbf{w}}.$
Since $\mathbf{y}_{G}$ and $\mathcal{E}_{\mathrm{ood}}$ are constants w.r.t. $G_{t}$ (detailed proof in Appendix A.1),
(7) $\nabla_{G_{t}}\log p_{t}\!\big(G_{t},\mathbf{y}_{G}\mid\mathcal{E}_{\mathrm{ood}}=\lambda\big)=\nabla_{G_{t}}\log p_{t}\!\big(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\mathrm{ood}}=\lambda\big).$
Thus the conditional reverse-time SDE becomes
(8) $\mathrm{d}G_{t}=\Big[\mathbf{f}_{t}(G_{t})-g_{t}^{2}\,\nabla_{G_{t}}\log p_{t}\!\big(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\mathrm{ood}}=\lambda\big)\Big]\mathrm{d}\bar{t}\;+\;g_{t}\,\mathrm{d}\overline{\mathbf{w}}.$
Score Decomposition. By Bayes’ rule,
(9) $\nabla_{G_{t}}\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)$
$=\nabla_{G_{t}}\log p_{t}\left(G_{t}\right)+\nabla_{G_{t}}\log p_{t}\left(\mathbf{y}_{G}\mid G_{t}\right)$
$+\nabla_{G_{t}}\log p_{t}\left(\mathcal{E}_{\text{ood }}=\lambda\mid G_{t},\mathbf{y}_{G}\right).$
where the last term, the OOD-ness score, is intractable to model directly. This is because modeling the conditional distribution,
$p_{t}\left(\mathcal{E}_{\text{ood }}=\lambda\mid G_{t},\mathbf{y}_{G}\right)$
would require an infeasible enumeration over a continuous OOD space and all corresponding graphs and labels.
A Tractable OOD Surrogate. From the causal view in §4.1, OOD graphs reside in low-density regions of the training joint $P_{\mathrm{tr}}(G,Y)$. We encode this by assigning higher OOD probability to lower joint density and posit the surrogate
(10) $p_{t}\left(\mathcal{E}_{\text{ood }}=\lambda\mid G_{t},\mathbf{y}_{G}\right)$ $\propto p_{t}\left(G_{t},\mathbf{y}_{G}\right)^{-\sqrt{\lambda}}$
$=\big[p_{t}(G_{t})\,p_{t}(\mathbf{y}_{G}\mid G_{t})\big]^{-\sqrt{\lambda}}.$
which yields
(11) $\nabla_{G_{t}}\log p_{t}\!\big(\mathcal{E}_{\mathrm{ood}}=\lambda\mid G_{t},\mathbf{y}_{G}\big)=$
$-\sqrt{\lambda}\,\Big(\nabla_{G_{t}}\log p_{t}(G_{t})+\nabla_{G_{t}}\log p_{t}(\mathbf{y}_{G}\mid G_{t})\Big).$
Plugging into (9) gives a simple and tractable conditional score:
(12) $\nabla_{G_{t}}\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)=$
$(1-\sqrt{\lambda})\Big(\nabla_{G_{t}}\log p_{t}(G_{t})+\nabla_{G_{t}}\log p_{t}(\mathbf{y}_{G}\mid G_{t})\Big).$
We use $0\leq\lambda\leq 1$ so that the guidance stays nonnegative and smoothly trades off between in-distribution sampling ($\lambda=0$) and exploration toward low-density regions ($\lambda\to 1$).
Learning Components. The tractable conditional score in Equation (12) necessitates estimating two key components: the unconditional score $\nabla_{G_{t}}\log p_{t}\left(G_{t}\right)$ and the class-conditional probability $p_{t}\left(\mathbf{y}_{G}\mid G_{t}\right)$. To this end, we first train a score network $\boldsymbol{s}_{\theta,t}\left(G_{t}\right)$ which takes the noisy graph $G_{t}$ and time $t$ as inputs to approximate the unconditional score $\nabla_{G_{t}}\log p_{t}\left(G_{t}\right)$. Concurrently, we train a time-dependent classifier $\phi_{t}(G_{t})$ to approximate the conditional class distribution $p_{t}\left(\mathbf{y}_{G}\mid G_{t}\right)$ using inputs $G_{t}$ and $t$. The classifier is trained on noisy pairs $(G_{t},\mathbf{y}_{G})$ to capture time-dependent feature dynamics, rather than on clean $(G_{0},\mathbf{y}_{G})$.
4.4. Working Principles of CODA
Substituting (12) into (8) defines the full sampler. The single hyperparameter $\lambda$ controls the amount of OOD exploration while retaining the class semantics via $\mathbf{y}_{G}$.
Assumptions (from §4.1). Each graph $G$ decomposes into $C$ (causal) and $S$ (non-causal/spurious), with $C\to Y$, $C\not\!\perp\!\!\!\perp S$, and $S\perp\!\!\!\perp Y\mid C$. In particular, $p_{t}(\mathbf{y}_{G}\mid G_{t})=p_{t}(\mathbf{y}_{G}\mid C_{t})$.
Proposition 4.1 (Causal structure preservation).
For $0\leq\lambda\leq 1$, the generation induced by (8)–(12) preserves the causal relation $C\!\to\!Y$ under the generated joint $\tilde{P}_{\lambda}(G,Y)$; equivalently, $\tilde{P}_{\lambda}(Y\mid G)=\tilde{P}_{\lambda}(Y\mid C)$ for all $\lambda\in[0,1]$.
Proof. Assumptions (cf. Section 4.1). Each graph decomposes as $G=(C,S)$ with $S\perp\!\!\!\perp Y\mid C$ and hence $P(Y\mid G)=P(Y\mid C)$. The time-dependent classifier is trained on noisy pairs so that $p_{t}(\mathbf{y}_{G}\mid G_{t})=p_{t}(\mathbf{y}_{G}\mid C_{t})$, implying $\nabla_{S_{t}}\log p_{t}(\mathbf{y}_{G}\mid G_{t})=0$.
(i) Vector-Field Structure. From (12),
(13) $\nabla_{G_{t}}\log p_{t}\!\big(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\mathrm{ood}}=\lambda\big)=$
$(1-\sqrt{\lambda})\Big(\underbrace{\nabla_{G_{t}}\log p_{t}(G_{t})}_{\text{label-free}}+\underbrace{\nabla_{G_{t}}\log p_{t}(\mathbf{y}_{G}\mid G_{t})}_{\text{$C$-only}}\Big).$
Projecting onto $S$-coordinates gives $\nabla_{S_{t}}\log p_{t}(\mathbf{y}_{G}\mid G_{t})=0$, so the label-guidance field has no component on $S$. Thus the $S$-dynamics are label-independent for all $\lambda$.
(ii) $\lambda=0$ anchors $C\!\to\!Y$. At $\lambda=0$, (12) reduces to standard class-conditional guidance $\nabla\log p_{t}(G_{t})+\nabla\log p_{t}(\mathbf{y}_{G}\mid G_{t})$ targeting $P_{\mathrm{tr}}(G\mid Y)$; consequently $\tilde{P}_{0}(\tilde{C}\mid\mathbf{y}_{G}=y)\approx P_{\mathrm{tr}}(C\mid y)$ and $\tilde{P}_{0}(Y\mid G)=\tilde{P}_{0}(Y\mid C)$.
(iii) Continuation in $\lambda$. For $0<\lambda\leq 1$, the overall field is scaled by $(1-\sqrt{\lambda})\geq 0$ and the label term remains $C$-only. Since the $S$-evolution is label-free, changing $\lambda$ alters how broadly we explore in $S$ but cannot introduce dependence of $Y$ on $S$ given $C$. Hence $P(Y\mid G)=P(Y\mid C)$ is preserved along the reverse-time flow for any $\lambda\in[0,1]$.
Therefore, for all $\lambda\in[0,1]$, $\tilde{P}_{\lambda}(Y\mid G)=\tilde{P}_{\lambda}(Y\mid C)$, i.e., the causal relation $C\to Y$ is preserved.
Proposition 4.2 (Spurious pattern reduction).
Let $\tilde{P}_{\lambda}$ be the generated joint induced by (12). For any fixed class $y$, write $\tilde{\pi}_{\lambda}(s\mid y):=\tilde{P}_{\lambda}(S=s\mid\mathbf{y}_{G}=y)$ and $\pi_{0}(s\mid y):=P_{\mathrm{tr}}(S=s\mid Y=y)$. Then for $0\leq\lambda_{1}<\lambda_{2}\leq 1$ and any measurable “top-mass” set $\mathcal{S}_{\tau}(y):=\{\,s:\pi_{0}(s\mid y)\geq\tau\,\}$, it holds that
$\tilde{\pi}_{\lambda_{2}}\big(\mathcal{S}_{\tau}(y)\mid y\big)\;\leq\;\tilde{\pi}_{\lambda_{1}}\big(\mathcal{S}_{\tau}(y)\mid y\big).$
In words: as $\lambda$ increases, the probability mass assigned to training-typical spurious patterns $S$ monotonically decreases under $\tilde{P}_{\lambda}$.
Proof. (i) Semantics are $C$-anchored. By Section 4.1, $S\perp\!\!\!\perp Y\mid C$, hence $P(Y\mid G)=P(Y\mid C)$. Moreover the time-dependent classifier is trained on noisy pairs so that $p_{t}(\mathbf{y}_{G}\mid G_{t})=p_{t}(\mathbf{y}_{G}\mid C_{t})$, implying $\nabla_{S_{t}}\log p_{t}(\mathbf{y}_{G}\mid G_{t})=0$. Thus the label-guidance term in (12) has no component on $S_{t}$.
(ii) Low-density reweighting induces power tempering. With the OOD surrogate $p_{t}(\mathcal{E}_{\mathrm{ood}}=\lambda\mid G_{t},\mathbf{y}_{G})\propto p_{t}(G_{t},\mathbf{y}_{G})^{-\sqrt{\lambda}}$, the class-conditional sampler satisfies
$p_{t}(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\mathrm{ood}}=\lambda)\;\propto\;p_{t}(G_{t}\mid\mathbf{y}_{G})^{\alpha(\lambda)},$
$\alpha(\lambda):=1-\sqrt{\lambda}\in(0,1].$
Marginalizing $C_{t}$ gives, for any fixed $y$,
$\tilde{\pi}_{\lambda}(s\mid y)\;\propto\;\int p_{t}(C_{t},s\mid y)^{\alpha(\lambda)}\,\mathrm{d}C_{t},$
followed by normalization over $s$. Note that $\alpha(\lambda)$ decreases as $\lambda$ increases.
(iii) Power tempering compresses high-probability advantages. Let $q_{s}:=\int p_{t}(C_{t},s\mid y)\,\mathrm{d}C_{t}=\pi_{0}(s\mid y)$. For any $s_{1},s_{2}$ with $q_{s_{1}}\geq q_{s_{2}}>0$ and any $\alpha\in(0,1]$,
$\frac{q_{s_{1}}^{\alpha}}{q_{s_{2}}^{\alpha}}=\Big(\frac{q_{s_{1}}}{q_{s_{2}}}\Big)^{\alpha}\leq\frac{q_{s_{1}}}{q_{s_{2}}}.$
Hence, as $\alpha(\lambda)$ decreases (i.e., $\lambda$ increases), the normalized mass shifts away from high-density (“typical”) $S$ configurations toward lower-density ones.
(iv) Conclusion. Combining (i)–(iii), increasing $\lambda$ cannot alter the $C$-anchored semantics, but it reduces the normalized probability of training-typical $S$ under $\tilde{\pi}_{\lambda}(\cdot\mid y)$, yielding the stated monotonicity on any top-mass set $\mathcal{S}_{\tau}(y)$.
5. Experiments
We first demonstrate the effectiveness of our diffusion models on graph OOD tasks. Then, we validate that the desired augmentation principles are successfully realized. Additional experiments regarding sensitivity of hyperparameter $\lambda$, as well as time and memory complexity, are presented in Appendix F and H.
5.1. Experimental Settings
Setup. We adopt the same evaluation metrics as in (gui2022good) for a fair comparison. The model that achieves the best performance on the OOD validation sets is then evaluated on the OOD test sets. Furthermore, to ensure fair comparison across all methods, we utilize the same GNN backbones—GIN (xu2018how) and GIN-Virtual (xu2018powerful; gilmer2017neural)—as applied in the GOOD benchmark (gui2022good) for each dataset. The experimental details, including evaluation metrics and hyperparameter configurations, are summarized in Appendix B and D.
Table 1. Performance on synthetic and real-world datasets. Bold numbers indicate the best performance, while the underlined numbers indicate the second best performance. Bootstrap test is used to assess significance with $20000$ resamples. The dagger symbol $\left({}^{\dagger}\right)$ indicates that our method significantly outperforms the best baseline with a p-value less than $0.05$.
Type Method Motif CMNIST Molhiv GOOD-SST2
base size color scaffold size length
General Generalization ERM $68.66\pm 4.25$ $51.74\pm 2.88$ $28.60\pm 1.87$ $69.58\pm 2.51$ $59.94\pm 2.37$ $81.30\pm 0.35$
IRM $70.65\pm 4.17$ $51.41\pm 3.78$ $27.83\pm 2.13$ $67.97\pm 1.84$ $59.00\pm 2.92$ $79.91\pm 1.97$
GroupDRO $68.24\pm 8.92$ $51.95\pm 5.86$ $29.07\pm 3.14$ $70.64\pm 2.57$ $58.98\pm 2.16$ $81.35\pm 0.54$
VREx $71.47\pm 6.69$ $52.67\pm 5.54$ $28.48\pm 2.87$ $70.77\pm 2.84$ $58.53\pm 2.88$ $80.64\pm 0.35$
DANN $65.47\pm 5.35$ $51.46\pm 3.41$ $29.14\pm 2.93$ $70.63\pm 1.82$ $\underline{62.38\pm 2.65}$ $79.71\pm 1.35$
Deep Coral $68.88\pm 3.61$ $53.71\pm 2.75$ $29.05\pm 2.19$ $68.61\pm 1.70$ $60.11\pm 3.53$ $79.81\pm 0.22$
Graph Generalization DIR $62.07\pm 8.75$ $52.27\pm 4.56$ $33.20\pm 6.17$ $68.07\pm 2.29$ $58.08\pm 2.31$ $77.65\pm 1.93$
GSAT $62.80\pm 11.41$ $53.20\pm 8.35$ $28.17\pm 1.26$ $68.66\pm 1.35$ $58.06\pm 1.98$ $81.49\pm 0.76$
CIGA $66.43\pm 11.31$ $49.14\pm 8.34$ $32.22\pm 2.67$ $69.40\pm 2.39$ $59.55\pm 2.56$ $80.44\pm 1.24$
Graph Augmentation DropNode $\underline{74.55\pm 5.56}$ $54.14\pm 3.11$ $33.01\pm 0.12$ $\underline{71.18\pm 1.16}$ $58.52\pm 0.49$ $81.14\pm 1.73$
DropEdge $45.08\pm 4.46$ $45.63\pm 4.61$ $22.65\pm 2.90$ $70.78\pm 1.38$ $58.53\pm 1.26$ $78.93\pm 1.34$
MaskFeature $64.98\pm 6.95$ $52.24\pm 3.75$ $44.85\pm 2.42$ $65.90\pm 3.68$ $62.30\pm 3.17$ $\underline{82.00\pm 0.73}$
FLAG $61.12\pm 5.39$ $51.66\pm 4.14$ $32.30\pm 2.69$ $68.45\pm 2.30$ $60.59\pm 2.95$ $77.05\pm 1.27$
M-Mixup $70.08\pm 3.82$ $51.48\pm 4.91$ $26.47\pm 3.45$ $68.88\pm 2.63$ $59.03\pm 3.11$ $80.88\pm 0.60$
G-Mixup $59.66\pm 7.03$ $52.81\pm 6.73$ $31.85\pm 5.82$ $70.01\pm 2.52$ $59.34\pm 2.43$ $80.28\pm 1.49$
GRATIN $72.45\pm 4.23$ $52.37\pm 3.17$ $32.77\pm 0.50$ $70.23\pm 2.29$ $60.28\pm 2.12$ $80.64\pm 0.51$
AIA $73.64\pm 5.15$ $\underline{55.85\pm 7.98}$ $\underline{36.37\pm 4.44}$ $71.15\pm 1.81$ $61.64\pm 3.37$ $81.69\pm 0.57$
CODA(Ours) $\mathbf{75.25\pm 3.84}^{\dagger}$ $\mathbf{60.81\pm 7.80}^{\dagger}$ $\mathbf{54.60\pm 2.27}^{\dagger}$ $\mathbf{72.67\pm 1.28}^{\dagger}$ $\mathbf{63.66\pm 1.20}^{\dagger}$ $\mathbf{82.69\pm 0.28}^{\dagger}$
Datasets. We evaluate on synthetic, semi-artificial, and real-world benchmarks from GOOD (gui2022good): GOOD-Motif, GOOD-CMNIST, GOOD-HIV, and GOOD-SST2. Following the official protocol, we use the base, size, color, scaffold, and length splits to induce diverse covariate shifts in structure, node features, and edge features. Although some datasets are not Web-native, the shifts emulate common Web scenarios: motif and size shifts mirror evolving community wiring and graph growth in social/hyperlink networks; color shifts stand in for feature/style changes (e.g., UI or content modality drift); scaffold splits mimic cold-start categories/items in e-commerce or knowledge graphs; and length shifts capture post/review length drift in social and content platforms. These settings therefore serve as Web-analogous stress tests for OOD generalization on graphs. Full dataset statistics, preprocessing details, and split constructions are provided in the Appendix D.
Baselines. We adopt 17 baselines, which can be divided into the following three specific categories:(i) general generalization algorithms, including ERM, IRM (arjovsky2019invariant), GroupDRO (sagawa2019distributionally), VREx (krueger2021out), DANN (ganin2016domain), Deep Coral (sun2016deep); (ii) graph generalization algorithms, including DIR (wu2022discovering), GSAT (miao2022interpretable), CIGA (chen2022learning), and (iii) graph data augmentation techniques, including DropNode (feng2020graph), DropEdge (rong2019dropedge), MaskFeature (thakoor2021large), FLAG (kong2022robust), M-Mixup (wang2021mixup), G-Mixup (han2022g), AIA (sui2024unleashing), GRATIN (abbahaddou2025graph).
5.2. Graph Out-of-Distribution Classification
The graph classification performances under covariate shift are presented in Table 1. As shown, CODA consistently outperforms all baseline methods across diverse covariate shifts and different datasets. On the synthetic dataset GOOD-Motif, CODA achieves a performance improvement of $6.59\%$ over ERM under base shift and $9.07\%$ under size shift. For the semi-artificial dataset GOOD-CMNIST, designed for node feature shifts, performance is significantly enhanced by $18.23\%$ compared to the leading graph augmentation method, AIA, and improved by $21.40\%$ over the best graph invariant learning method, DIR. In the real-world molecular dataset GOOD-HIV, where covariate shifts occur in graph structure, node features, and edge features simultaneously, CODA outperforms ERM by $3.09\%$ on scaffold shift and by $3.72\%$ on size shift. For the real-world sentiment analysis dataset GOOD-SST2, while AIA is outperformed by MaskFeature by $0.31\%$, CODA exceeds MaskFeature by $0.69\%$.
We assess statistical significance using a paired nonparametric bootstrap test (davison1997bootstrap) at the $5\%$ level to compare CODA with leading baselines; details are in Appendix G. Overall, these results demonstrate that no baseline graph invariant learning or graph data augmentation methods consistently outperform each other under various covariate shifts. CODA enhances environmental exploration by generating OOD graphs in a controlled manner while preserving stable features. Consequently, CODA reliably improves performance across different datasets facing various covariate shifts.
5.3. OOD Controlled Graph Generation
In this section, we present both qualitative and quantitative experiments to demonstrate the effectiveness of our OOD diffusion augmentation framework. The experiments are conducted using the GOOD-Motif-base, GOOD-SST2-length and GOOD-HIV-scaffold datasets. Additionally, we present a comparative analysis of CODA against various augmentation methods, evaluating their respective capacities to explore environments while preserving label-determining stable structures. We further provides visualizations of the OOD graphs generated by CODA for the GOOD-Motif-basis and GOOD-Molhiv-scaffold datasets.
Training Distribution Augmentation with Control. We first validate that our framework can explore the space of the original distribution $P_{\mathrm{tr}}(G,Y)$ and generate an augmented distribution $\tilde{P}_{\mathrm{tr}}(G,Y)$ in a controlled manner. To quantify deviation from training graphs, we adopt the GIN-based metrics of (thompson2022on), which are expressive and computationally efficient for graph generators. Following their recommendation, we report Maximum Mean Discrepancy (MMD) with an RBF kernel as a robust measure of distributional shift and diversity. We vary $\lambda\in[0,1)$ in steps of $0.1$ to generate ten augmented sets (each the same size as the training set) and compute MMD (RBF) between each augmented set and the original training set. Estimation details are in the Appendix.
GOOD-Motif-basis. Figure 2(a) shows that MMD (RBF) between $P_{\mathrm{tr}}(G,Y)$ and $\tilde{P}_{\mathrm{tr}}(G,Y)$ increases monotonically with $\lambda$. At $\lambda=0.0$, MMD is $0.072\pm 0.002$; by $\lambda=0.9$, it is approximately $1.5\times$ larger. Thus, CODA generates OOD graphs in a tunable manner by adjusting $\lambda$, while also increasing diversity as the exploration strength grows.
GOOD-SST2-length. Here, $\lambda$ controls OOD size (e.g., $\lambda=0.1$ adds one unit to graph size relative to training). We again sweep $\lambda$ in $[0,1)$ by $0.1$ to obtain ten augmented sets. As shown in Figure 2(b), MMD (RBF) rises with $\lambda$: from $0.032\pm 0.005$ at $\lambda=0.0$ to roughly $25\times$ that value at $\lambda=0.9$.
Stable Pattern Preservation. We use a pretrained graph transformer $\phi_{t}$, trained on noisy graphs from $P_{\mathrm{tr}}$, as a proxy to assess whether augmented graphs retain label-determining stable patterns. Specifically, we compute $p(\mathbf{y}_{G}\mid\tilde{G}_{0})$ and report $\mathbb{E}_{\tilde{G}_{0}\sim\tilde{P}}[p(\mathbf{y}_{G}\mid\tilde{G}_{0})]$. On GOOD-Motif-basis (Figure 2(c)), this expectation remains above $0.95$ across $\lambda$; on GOOD-SST2-length (Figure 2(d)), it exceeds $0.94$. These results indicate that CODA generates OOD graphs while preserving stable, label-determining structure.
Comparison with augmentation baselines. Figure 2 compares methods on (i) environmental exploration and (ii) preservation of stable structures. G-Mixup mixes continuous labels, making discrete structure preservation ill-posed; we therefore omit that metric for G-Mixup. Chemistry-specific metrics (e.g., FCD (preuer2018frechet), RDKit-based scores (landrum2016rdkit)) require valid discrete molecules and are incompatible with baselines that output continuous graph features (e.g., AIA), so we focus on general graph metrics. Across GOOD-Motif-basis and GOOD-SST2, DropEdge explores more aggressively than G-Mixup and AIA but often distorts stable patterns. In contrast, CODA achieves stronger, tunable exploration via $\lambda$ (Figure 2(a) and Figure 2(b)) while consistently preserving stable structures better than the baselines (Figure 2(c) and Figure 2(d)).
Figure 2. Augmentation comparison on GOOD-Motif-basis and GOOD-SST2-length. (a) GOOD-Motif-basis: distribution distance between $P_{\mathrm{tr}}(G,Y)$ and $\tilde{P}_{\mathrm{tr}}(G,Y)$ (MMD-RBF). (b) GOOD-SST2-length: distribution distance (MMD-RBF). (c) GOOD-Motif-basis: stable pattern preservation probability (probability augmented graphs preserve label-determining stable patterns). (d) GOOD-SST2-length: stable pattern preservation probability.
Figure 3. (Left): Distance between the original GOOD-HIV graph distribution and the augmented graph distribution. (Right): The validity of the OOD molecules and the expected probabilities that the OOD GOOD-HIV molecules retain stable patterns.
[…]
Ablation study. We present experiments to evaluate the impact of exploration guidance ($\lambda$) and stable patterns preservation guidance ($\alpha$). The results are summarized in Table 2. As shown, incorporating $\alpha$ guidance improves the OOD performance of the diffusion model trained on unlabeled graphs by ensuring that the generated graphs retain the stable patterns that determine the labels. Without $\alpha$ guidance, a diffusion model guided only by $\lambda$ tends to push the augmented graphs into arbitrary OOD regions, which negatively impacts performance to some extent. Ultimately, the combination of both $\alpha$ and $\lambda$ guidance enables the augmented distribution to capture both stable patterns and novel environmental patterns, resulting in the best overall performance.
Table 2. Performance of CODA w/o environmental exploration guidance and stable pattern preservation guidance on synthetic and real-world datasets. Bold numbers indicate the best performance.
$\lambda$ $\alpha$ Motif-base Molhiv-scaffold GOOD-SST2-length
ERM $68.66\pm 4.25$ $69.58\pm 2.51$ $81.30\pm 0.35$
$68.55\pm 6.04$ $68.94\pm 1.26$ $80.86\pm 0.76$
$66.25\pm 7.42$ $70.01\pm 1.71$ $78.87\pm 3.04$
$74.57\pm 4.50$ $71.71\pm 1.77$ $81.59\pm 0.65$
CODA $\mathbf{75.25\pm 3.84}$ $\mathbf{72.67\pm 1.28}$ $\mathbf{82.69\pm 0.28}$
Visualization of OOD graphs We further demonstrate the efficacy of CODA by visualizing the OOD graphs generated by our approach in Table 3. In this visualization, three label-determining motifs—house, cycle, and crane—are highlighted in red, while the three environmental base graphs in the training distribution—wheel, tree, and ladder—are indicated in green. As illustrated in Table 3, increasing $\lambda$ leads to gradual modifications in the structures of the base graphs, while the motifs remain preserved.
Table 3. Visualizations of the augmented GOOD-Motif-base graphs generated by CODA. The graphs $\tilde{G}$ with $\lambda=0.1$, $\lambda=0.2$, and $\lambda=0.3$ represent the OOD graphs generated by CODA under different values of $\lambda$.
Original Graphs $\tilde{G}(\lambda=0.1)$ $\tilde{G}(\lambda=0.2)$ $\tilde{G}(\lambda=0.3)$
House Class
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Cycle Class
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Crane Class
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Visualization of OOD molecules We showcase the effectiveness of CODA by visualizing the OOD molecules generated by our approach in Table 4. As illustrated in Table 4, increasing $\lambda$ results in gradual modifications to the structures of the molecular graphs, while the properties of the molecules remain preserved. This controlled exploration allows CODA to produce chemically meaningful and diverse OOD molecules without disrupting their underlying validity. Notably, the changes remain smooth and interpretable across different $\lambda$ values, highlighting the tunable nature of our generation process.
Table 4. Visualizations of the augmented GOOD-Molhiv-scaffold graphs generated by CODA. The graphs $\tilde{G}$ with $\lambda=0.1$, $\lambda=0.2$, and $\lambda=0.3$ represent the OOD graphs generated by CODA under different values of $\lambda$.
Original Graphs $\tilde{G}(\lambda=0.1)$ $\tilde{G}(\lambda=0.2)$ $\tilde{G}(\lambda=0.3)$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
6. Conclusion
We presented CODA, a score-based out-of-distribution graph generation framework for addressing covariate shift in graph learning. By coupling a conditional score with an exploration parameter, CODA explores low-density regions beyond the training support in a controlled manner while preserving label-determining stable patterns. Unlike edit-based augmentations, our approach jointly synthesizes structures, node features, and edge features and does not require explicit decomposition into stable and environmental components. Across synthetic, semi-artificial, and real benchmarks from GOOD, CODA consistently improves OOD generalization over invariant-learning and augmentation baselines. The exploration knob $\lambda$ yields monotonic increases in MMD while maintaining a high probability of retaining stable patterns, demonstrating both controllability and fidelity. Given that distribution shift is pervasive on the Web (e.g., evolving social, hyperlink, and interaction graphs), these results indicate practical value for robust web-scale modeling.
Overall, CODA provides a principled and controllable path to web-relevant graph augmentation, improving robustness to the distribution shifts that dominate modern web applications.
Appendix A Theory
A.1. Deriving the conditional reverse-time SDE
[…]
Proof:
$\log p_{t}\left(G_{t},\mathbf{y}_{G}\mid\mathcal{E}_{\text{ood }}=\lambda\right)=$ $\log p_{t}\left(G_{t},\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)$
- $\log p_{t}\left(\mathcal{E}_{\text{ood }}=\lambda\right)$
Since $p_{t}\left(\mathcal{E}_{\text{ood }}=\lambda\right)$ is independent of $G_{t}$, $\nabla_{G_{t}}\log p_{t}\left(\mathcal{E}_{\text{ood }}=\lambda\right)=0$. Therefore,
[…]
Additionally,
$\log p_{t}\left(G_{t}\mid\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)$ $=\log p_{t}\left(G_{t},\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)$
$-\log p_{t}\left(\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)$
Since $p_{t}\left(\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)$ is independent of $G_{t}$,
$\nabla_{G_{t}}\log p_{t}\left(\mathbf{y}_{G},\mathcal{E}_{\text{ood }}=\lambda\right)=0$
Therefore,
[…]
Finally,
[…]
Appendix B Metrics for measuring distributional differences
In this section, we provide detailed implementation steps for measuring the distributional differences between the augmented dataset and the training dataset. Following (thompson2022on), we first use an untrained random GIN, $h$, to extract graph embeddings from both the augmentation distribution and the training distribution. The maximum mean discrepancy (MMD) is then computed to quantify the dissimilarity between the graph embedding distributions:
$)]+\displaystyle\operatorname{MMD}^{2}(P\|\tilde{P})=\mathbb{E}_{g,\tilde{g}\sim P}[k(h(g),h(\tilde{g}$))]+
$)]\displaystyle\mathbb{E}_{g,\tilde{g}\sim\tilde{P}}[k(h(g),h(\tilde{g}))]-2\mathbb{E}_{g\sim P,\tilde{g}\sim\tilde{P}}[k(h(g),h(\tilde{g}$))]
where $k(\cdot,\cdot)$ is the RBF kernel proposed by (you2018graphrnn). As recommended by (thompson2022on), the MMD RBF scalar is also one of the most reliable metrics for measuring distributional differences:
$k\left(h(g),h(\tilde{g})\right)=\exp\left(-d\left(h(g),h(\tilde{g})\right)/2\sigma^{2}\right)$
Additionally, we employ the Earth Mover’s Distance (EMD) from (thompson2022on) to compute pairwise distances $d(\cdot,\cdot)$.
Appendix C Implementations of CODA.
Directly applying this framework to graphs is inadequate for addressing complex covariate shifts, especially those that simultaneously affect feature and structural distributions. To overcome this, we explicitly model the joint diffusion of node features and adjacency matrices, representing the perturbed graph as $\left\{G_{t}=\left(\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)\right\}_{t=0}^{T}$ using a set of SDEs for Equation 8:
(14) $\left\{\begin{aligned} \mathrm{d}\boldsymbol{X}_{t}=&{\left[\mathbf{f}_{1,t}\left(\boldsymbol{X}_{t}\right)-(1-\sqrt{\lambda})g_{1,t}^{2}\nabla_{\boldsymbol{X}_{t}}\log p_{t}\left(\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)\right.}\\ &\left.-(1-\sqrt{\lambda})g_{1,t}^{2}\nabla_{\boldsymbol{X}_{t}}\log p_{t}\left(\mathbf{y}_{G}\mid\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)\right]\mathrm{d}\bar{t}\\ &+g_{1,t}\mathrm{~d}\overline{\mathbf{w}}_{1}\\ \mathrm{~d}\boldsymbol{A}_{t}=&{\left[\mathbf{f}_{2,t}\left(\boldsymbol{A}_{t}\right)-(1-\sqrt{\lambda})g_{2,t}^{2}\nabla_{\boldsymbol{A}_{t}}\log p_{t}\left(\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)\right.}\\ &\left.-(1-\sqrt{\lambda})g_{2,t}^{2}\nabla_{\boldsymbol{A}_{t}}\log p_{t}\left(\mathbf{y}_{G}\mid\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)\right]\mathrm{d}\bar{t}\\ &+g_{2,t}\mathrm{~d}\overline{\mathbf{w}}_{2}.\end{aligned}\right$.
where $\mathbf{f}_{t}(\boldsymbol{X},\boldsymbol{A})=\left(\mathbf{f}_{1,t}(\boldsymbol{X}),\mathbf{f}_{2,t}(\boldsymbol{A})\right)$ and $g_{t}=\left(g_{1,t},g_{2,t}\right)$ representing the drift and diffusion coefficients, respectively. The reverse-time processes are captured by standard Wiener processes $\overline{\mathbf{w}}_{1}$ and $\overline{\mathbf{w}}_{2}$, with $\mathrm{d}\bar{t}$ indicating an infinitesimally small negative time step. We train one graph transformer (dwivedi2020generalization; vignac2022digress), denoted as $s_{\theta,t}=\left(s_{\theta_{1},t},s_{\theta_{2},t}\right)$ to closely estimate the partial score functions:
$s_{\theta,t}=\big(s_{\theta_{1},t},\,s_{\theta_{2},t}\big)\approx\Big(\nabla_{\boldsymbol{X}_{t}}\log p_{t}(\boldsymbol{X}_{t},\boldsymbol{A}_{t}),\;\nabla_{\boldsymbol{A}_{t}}\log p_{t}(\boldsymbol{X}_{t},\boldsymbol{A}_{t})\Big).$
Therefore, the denoising graph transformer is only trained with unlabelled graphs without $\lambda$ values and graph labels. The graph transformer takes perturbed node features $X_{t}$, adjacency matrices $A_{t}$ and normalized timestep as input. The timestep $t$ value is treated as a global graph feature, and an embedding layer is used to embed $t$.
We also use a graph transformer model $\phi_{t}$ with the same architecture to predict the class label of the noisy graphs $G_{t}=\left(\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)$ at time step $t$. The target class $j$ probability $p_{t}\left(\mathbf{y}_{G}=j\mid\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)$ is then given by:
$p_{t}\left(\mathbf{y}_{G}=j\mid\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)=\frac{e^{\phi_{t}\left(\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)_{[j]}}}{\sum_{j=1}^{M}e^{\phi_{t}\left(\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)_{[j]}}}.$
Once both the score transformer and the classifier are trained, we use them to compute the conditional partial scores in Equation (8) during the sampling process. It is noteworthy that while our generation process leverages the representational power of graph transformers, subsequent evaluations for graph OOD classification rely on augmenting only simple GNN backbones with these generated samples (as detailed in Section 5.1).
We adopt the two popular time-dependent hyperparameters $\alpha_{1,t}$ and $\alpha_{2,t}$ for the target class probability predicted by $\phi_{t}$. These hyperparameters are defined as follows:
[…]
where $\alpha_{t}=\left(\alpha_{1,t},\alpha_{2,t}\right)$, $r_{1}$ and $r_{2}$ are the weights for node features and adjacency matrices respectively, and $\|\cdot\|$ is the entry-wise matrix norm.
Intuitively, at the early stages of the reverse-time SDEs, the graphs are highly perturbed, resembling the prior noise distribution. Therefore, the classifier cannot accurately approximate the target class probability. Consequently, in the initial denoising steps, we focus more on guiding the reverse-time SDEs towards the low-density (i.e. OOD distribution) regions. As the reverse-time SDEs progressively denoise the graphs, we introduce guidance to direct the reverse-time SDEs towards regions exhibiting the desired stable patterns and the explored OOD environmental patterns.
Appendix D Experimental Details
D.1. Dataset Details
We utilize six datasets from the GOOD benchmark (gui2022good), including GOOD-Motif-base, GOOD-Motif-size, GOOD-CMNIST-color, GOOD-HIV-scaffold, GOOD-HIV-size, and GOOD-SST2-length. The GOOD benchmark (gui2022good) is the state-of-the-art framework for systematically evaluating graph OOD generalization. It carefully designs data environments to induce reliable and valid distribution shifts. The selected datasets span a diverse range of domains, covering covariate shifts in general graphs, image-transformed graphs, molecular graphs, and natural language sentiment analysis graphs. The dataset details are as follows:
  • GOOD-Motif: GOOD-Motif is a synthetic dataset from Spurious-Motif (wu2022discovering) specifically designed to investigate structure shifts. Each graph consists of an environmental base graph connected to a label-determining motif. The two primary covariate shift domains are the base graph type and graph size. For base covariate shift, the training distribution includes graphs with wheel, tree, and ladder base structures, while the validation set features star base graphs, and the test set contains path base graphs. For size covariate shift, the training distribution consists of graphs with sizes ranging from $6$ to $45$ nodes, the validation set contains graphs with sizes between $20$ and $75$ nodes, and the test set comprises graphs with sizes ranging from $68$ to $155$ nodes.
  • GOOD-CMNIST: GOOD-CMNIST is a semi-synthetic dataset designed to investigate node feature shifts. It consists of graphs transformed from MNIST handwritten digit images using superpixel techniques (monti2017geometric). Node color features are manually applied, making the color shift environment independent of the underlying structure. Specifically, for covariate shift, digits are colored using seven different colors. The training distribution includes digits colored with the first five colors, while the validation and test distributions contain digits with the remaining two colors, respectively.
  • GOOD-HIV: GOOD-HIV is a small-scale, real-world molecular dataset sourced from MoleculeNet (wu2018moleculenet). The nodes in these molecular graphs represent atoms, and the edges represent chemical bonds. This dataset is designed to study node feature shifts, edge feature shifts, and structure shifts. The two covariate shift domains are scaffold graph type and molecular graph size. For the scaffold covariate shift, environments are partitioned based on the Bemis-Murcko scaffold (bemis1996properties), a two-dimensional structural base that does not determine a molecule’s ability to inhibit HIV replication. For the size covariate shift, the training distribution consists of molecular graphs ranging in size from $17$ to $222$ atoms. The validation set contains molecules with sizes between $15$ and $16$ atoms, while the test set includes molecules with sizes from $2$ to $14$ atoms.
  • GOOD-SST2: GOOD-SST2 is a real-world natural language sentimental analysis dataset from (yuan2022explainability), designed to investigate node feature shifts and structure shifts. Each graph is derived from a sentence, transformed into a grammar tree, where nodes represent words, and node features are corresponding word embeddings. The task is to predict the sentiment polarity of each sentence. Sentence length is chosen as the covariate shift environment, as sentence length should not inherently affect sentiment polarity. For the length covariate shift, the training distribution consists of grammar graphs with sizes ranging from $1$ to $7$ nodes, the validation distribution includes graphs with sizes from $8$ to $14$ nodes, and the test distribution contains graphs with sizes from $15$ to $56$ nodes.
D.2. Implementation settings
Diffusion models: Following (jo2022score), we preprocess each graph into two matrices: $\boldsymbol{X}\in\mathbb{R}^{n\times a}$ for node features, and $\boldsymbol{A}\in\mathbb{R}^{n\times n\times b}$ for adjacency and edge features. Here, $n$ represents the maximum number of nodes in a graph for the given dataset, while $a$ and $b$ denote the dimensions of node features and edge features, respectively. The graph structure, including edge features, is encoded in $\boldsymbol{A}$. For the GOOD-Motif dataset, $a$ corresponds to the node degree of a node. In GOOD-CMNIST, each node feature is the concatenation of its degree and color. In GOOD-SST2, the node feature is the word embedding. In the molecular dataset GOOD-HIV, $a$ represents possible atom types and $b$ denotes the types of bonds (e.g., single, double, triple). All molecules are converted to their kekulized form, with hydrogens removed using the RDKit library (landrum2016rdkit). Additionally, we apply the valency correction proposed by (zang2020moflow) to post-process the generated molecules.
We train a graph transformer model (dwivedi2020generalization; vignac2022digress), $s_{\theta,t}$, to approximate the partial score functions for the unlabelled graphs in the OOD training set and evaluate them on the OOD validation set. In line with (jo2022score), we use VP or VE SDEs to model the diffusion process for both node features and adjacency matrices. The specific details of the diffusion models are provided in Table 5.
We also train a graph transformer model, $\phi_{t}$, with the same architecture described in Table 5, to predict the class labels of the noisy graphs $G_{t}=\left(\boldsymbol{X}_{t},\boldsymbol{A}_{t}\right)$ at each time step $t$.
Table 5. Hyperparameters of diffusion models.
Hyperparameter Motif CMNIST Molhiv GOOD-SST2
$s_{\theta}$ Number of graph transformer layers 88 88 99 88
Number of attention heads 88 88 88 88
Hidden dimension of $\boldsymbol{X}$ $256$ $256$ $256$ $256$
Hidden dimension of $\boldsymbol{A}$ $64$ $64$ $64$ $64$
SDE for $\boldsymbol{X}$ Type VP VP VP VP
Number of sampling steps 1000 1000 1000 1000
$\beta_{\text{min }}$ 0.1 0.1 0.1 0.1
$\beta_{\max}$ 1.0 1.0 1.0 1.0
SDE for $\boldsymbol{A}$ Type VP VP VE VP
Number of sampling steps 1000 1000 1000 1000
$\beta_{\text{min }}$ 0.1 0.1 0.2 0.2
$\beta_{\max}$ 1.0 1.0 1.0 0.8
Solver Type EM + Langevin EM + Langevin Reverse EM
SNR $0.2$ $0.2$ $0.0$ $0.0$
Scale coefficient $0.7$ $0.7$ $0.0$ $0.0$
Train Optimizer AdamW AdamW AdamW AdamW
Learning rate $4\times 10^{-4}$ $4\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$
Weight decay $\times 10^{-12} 1×1$$\times 10^{-12}$ $\times 10^{-12} 1×1$$\times 10^{-12}$
Batch size $128$ $64$ $512$ $64$
EMA $0.999$ $0.999$ $0.999$ $0.999$
Graph Out-of-Distribution Classification: Following prior work (gui2022good; ligraph), we employ GIN-Virtual (xu2018powerful; gilmer2017neural) as the GNN backbone for the GOOD-CMNIST, GOOD-HIV, and GOOD-SST2 datasets. For the GOOD-Motif dataset, we adopt GIN (xu2018how). To ensure a fair comparison across all methods, we utilize the same GNN backbone architecture for all models.
For each experiment, we select the best checkpoints for OOD testing based on the performance on the OOD validation sets. All experiments are optimized using the Adam optimizer, with weight decay selected from $\{0,1\times 10^{-2},1\times 10^{-3},1\times 10^{-4}\}$ and a dropout rate of $0.5$. The number of convolutional layers in the GNN models is tuned from the set $\{3,5\}$, with mean global pooling and ReLU activation. The hidden layer dimension is set to $300$. We explore the maximum number of epochs from $\{100,200,500\}$, the initial learning rate from $\{1\times 10^{-3},3\times 10^{-3},5\times 10^{-3},1\times 10^{-4}\}$, and the batch size from $\{32,64,128\}$. All models are trained to convergence.
For computation, we typically run each experiment on an NVIDIA GeForce RTX 4090. We report results as the mean and standard deviation across 10 random runs for all experiments.
We perform a grid search for the hyperparameter $\alpha\in\{0.1,0.5,1.0\}$ across all datasets. For $\lambda$, the grid search is tailored to each dataset. Specifically, we explore $\lambda\in\{0.01,0.02,0.03,0.04,0.05\}$ for the GOOD-Motif-base dataset and $\lambda\in\{0.1,0.2,0.3\}$ for GOOD-HIV-scaffold dataset. For GOOD-CMNIST-color, we tune $\lambda\in\{0.05,0.1\}$. In the case of GOOD-SST2-length, where $\lambda=0.01$ corresponds to an increase of one node in the graph size relative to the training distribution, we expand the grid search to $\lambda\in\{0.01,0.02,...,0.14\}$. Similarly, for GOOD-Motif-size, where $\lambda=0.01$ reflects an increase of one node, we use a search space of $\lambda\in\{0.01,0.02,0.03,0.04,0.05\}$. For GOOD-HIV-size, where $\lambda=0.01$ corresponds to a decrease of ten nodes in graph size from the training distribution, we also use $\lambda\in\{0.01,0.02,0.03,0.04,0.05\}$. Since this hyperparameter tuning is performed during the sampling phase rather than the training phase, it is not computationally intensive.
Appendix E Baseline settings
The implementation details for GNN backbones and hyperparameter tuning are consistent with those outlined in Appendix D.2. For methods including ERM, IRM (arjovsky2019invariant), GroupDRO (sagawa2019distributionally), VREx (krueger2021out), DANN (ganin2016domain), Deep Coral (sun2016deep), DIR (wu2022discovering), DropNode (feng2020graph), DropEdge (rong2019dropedge), MaskFeature (thakoor2021large), FLAG (kong2022robust), M-Mixup (wang2021mixup), and G-Mixup (han2022g), we report results from the study in (ligraph), which uses the same GNN backbones and hyperparameter tuning as specified in Appendix D.2. For GSAT (miao2022interpretable), CIGA (chen2022learning), and AIA (sui2024unleashing), we use their publicly available source code, adopting default settings and hyperparameters as detailed in their papers.
Appendix F Sensitivity of hyperparameter $\lambda$
We determine the value of $\lambda$ by evaluating its effectiveness on OOD validation sets across various datasets. The sensitivity of our method to different $\lambda$ values is illustrated in Tables 6, 7 and 8.
Table 6. Performance on the GOOD-Motif-base dataset across varying $\lambda$ values.
$\lambda$ GOOD-Motif-base
0.01 91.80
0.02 92.97
0.03 92.83
0.04 93.03
0.05 92.77
Table 7. Performance on the GOOD-HIV-scaffold dataset across varying $\lambda$ values.
$\lambda$ GOOD-HIV-scaffold
0.1 78.94
0.2 77.73
0.3 77.42
Table 8. Performance on the GOOD-CMNIST-color dataset across varying $\lambda$ values.
$\lambda$ GOOD-CMNIST-color
0.05 68.66
0.1 67.91
Appendix G Statistical Significance
Statistical significance is evaluated via a paired nonparametric bootstrap test (davison1997bootstrap) at the $5\%$ level. The observed gains are significant across settings: on GOOD-Motif (base), CODA improves over DropNode by $0.70\%$ with $p=0.0305$; on GOOD-Motif (size), over AIA by $4.96\%$ with $p=$ 0.0015 ; on GOOD-CMNIST (color), over AIA by $18.23\%$ with $p=4.9998\times 10^{-5}$; on GOOD-HIV (scaffold), over DropNode by $1.49\%$ with $p=0.0147$; on GOOD-HIV (size), over DANN by $1.28\%$ with $p=0.0062$; and on GOOD-SST2 (length), over MaskFeature by $0.69\%$ with $p=0.0207$.
Appendix H Time and memory complexity
Our pipeline consists of two stages: data augmentation, and trainng of the GNN classifier on augmented graphs. The second stage of classification follows general GNN training setup, without introducing additional complexity: the time/memory complexity per layer of using Graph Isomorphism Networks (GIN) as backbone is $\Theta(n+e)$, where $n$ is the number of nodes and $e$ is the number of edges. For the data augmentation stage, we introduce graph transformer whose memory and time complexity per layer is $\Theta(n^{2})$. This arises from the computation of attention scores and predictions across each edge.